1. Chef wants to diversify with a new line of cooking utensils. The project costs $25 million and generates earnings before interest and taxes (EBIT) of $3,600,000 every year forever. The unlevered cost of capital is 11%. They will issue $15 million in debt with an interest rate of 8%. The debt is never repaid. The tax rate is 30%. The debt ratio (debt to value) is 60%. Use flow-to-equity (FTE) to find the project’s value.

Unless stated otherwise, compounding is annual and payments occur at the end of the period.
Answer #1
1] Value of levered equity = 11%+(11%-8%)*(1-30%)*0.6/0.4 = 14.15%
2] EBIT $       36,00,000
Less: Interest on pepetual debt = 15000000*8% = $       12,00,000
EBT $       24,00,000
Tax at 30% $         7,20,000
NI $       16,80,000
3] Project's value = 1680000/14.15% = $   1,18,72,792